Circular No. 6673

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997cp IN ANONYMOUS GALAXY The Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC

6639) reports the discovery of an apparent supernova (R = 20.3) on V and R CCD images taken on June 2 by S. Sabine at the Mount Stromlo 1.27-m telescope (+ Macho Camera). SN 1997cp is located near Abell 3806 at α = 21^{h}44^{m}36^{s}.50, δ = −57^{o}25′30′′.4 (equinox 2000.0). A nearby star has position end figures 36^{s}.77, 31′′.9. SN 1997cp was also present on similar frames taken on June 3, but it is not visible on images taken on May 11 (limiting mag R = 21.2). Finding charts are available at

http://msowww.anu.edu.au/~reiss/Abell SNSearch/.

CYGNUS X-3 M. L. McCollough for the Compton Gamma Ray Observatory BATSE

team; E. B. Waltman, Naval Research Laboratory; and G. G. Pooley, Mullard Radio Astronomy Observatory, Cambridge, report: “Very low flux levels have been observed in radio and hard x-rays from Cyg X-3 since May 29. Very low radio fluxes of < 10 mJy have been observed at 2 and 8 GHz with the Green Bank Interferometer and at 15 GHz with the Ryle Telescope. Daily averages for May 29 to June 1 of the hard-x-ray flux in the band 20–100 keV have been below the BATSE one day detection limit of 100 mCrab. In Cyg X-3, periods of very low radio flux have durations from days to a few weeks and are followed by major radio flares (5 Jy or larger). A period of very low radio emission in 1991 coincided with the observation of strong infrared emission lines in Cyg X-3. Observations at other wavelengths, especially infrared, are encouraged.”

SUPERNOVA 1997bs IN NGC 3627 Unfiltered CCD magnitude estimates by S. Moretti and S. Tomaselli,

Forlí, Italy: May 2.92 UT, 17.5; 27.88, 18.6 (GSC comparison stars).

COMET 29P/SCHWASSMANN-WACHMANN 1 Total magnitude estimates, visual unless otherwise noted: May 1.95

UT, 13.1 (J. Carvajal, Teruel, Spain, 0.44-m reflector); 30.17, 11.4 (P. E. Roques, Williams, AZ, 0.25-m reflector + CCD + Kron-Cousins V filter); June 1.90, 11.0 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 3.19, 12.7 (A. Hale, Cloudcroft, NM, 0.41-m reflector).

1997 June 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green